A projectile is fired upward from the ground with an initial velocity of 571 feet per second. Assume that the height of the projectile at any time t can be desceibed by the polynomial function P(t)=-16t2(squared) +571t. Find the height of the projectile at each given time.
3 seconds?
6 seconds?
10 seconds?
12 seconds?
Explain why the height changes as time passes.

Answer by Alan3354(21608)

A projectile is fired upward from the ground with an initial velocity of 571 feet per second. Assume that the height of the projectile at any time t can be desceibed by the polynomial function P(t)=-16t2(squared) +571t. Find the height of the projectile at each given time.
P(t) = -16t^2 + 571t
3 seconds?
P(3) = -144 + 1713 = 1569 feet
6 seconds?
P(6) = -576 + 3426 = 2850 feet
10 seconds?
P(10) = -1600 + 5710 = 4110 feet
12 seconds?
P(12) = -2304 + 6852 = 4548 feet
Explain why the height changes as time passes.
Because it's moving. Jeez